OpenStudy (anonymous):

Two trains left the station at the same time, one heading due east at 40 miles per hour and the other heading due west at 50 miles per hour. How far apart were they after 4 hours? 160 miles 200 miles 360 miles 400 miles

OpenStudy (taylor0402):

The speeds are added when going in opposite directions d = r*t

OpenStudy (taylor0402):

so 4*40=160 50*4=200. since it was opposite directions you add. so 160+200=???

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (taylor0402):

no problemo